OpenStudy (anonymous):

the distance from cambridge to ottawa is 480 km. gerry is able to drive at 100km/h on highway 401, but can only average 60km/h on highway 7. if the trip takes 6h, how far did he travel on highway 401?

OpenStudy (anonymous):

x = 100t where x = distance travelled on 401, t = time on 401 480-x = 60(t-6) solve for x and t

OpenStudy (anonymous):

sorry second equation is wrong it should be 480-x = 60(6-t)

OpenStudy (anonymous):

substitute x=100t in the above 480 - 100t = 360 - 60t 120 = 40t t = 3 hours distance travelled on highwat 401 = 3 * 100 = 300 km